John wants to hire an editor to check his thesis paper. The editor charges a flat rate of $50 and $2 per page. Write a polynomial to describe the cost. Find the cost of editing 75 pages.
Calculations:

Answer: Polynomial: ____________________Cost:_________________

Answer by harpazo(655)

John wants to hire an editor to check his thesis paper. The editor charges a flat rate of $50 and $2 per page. Write a polynomial to describe the cost. Find the cost of editing 75 pages.
Let C = cost for editing 75 pages
Let x = number of pages
C = 50 + 2x
C = 50 + 2(75)
C = 50 + 150
C = $200